Tech Staff Meeting Summary 4/30/02

Current Top Priorities are:Complete DBB, ANSS

CVTN is running. Need one more visit to get comm going.

CBHT comm. Switch combined with dramatic k2ew config changes appeared to work. Steve said a new firmware upgrade may also solve the problem.

HDBT. Need a fence and gravel for ground cover.

NAIT. NWS does not know when the VPN will be installed. Mitch will ask the NWS Houston contact for more info.

CUET. Provided additional information to Pat Bohannan at TEMA regarding the software connections. He's working on the firewall issues (we hope).

SFTN. K2 and MSS100 are fine. No comm out the ethernet. Needs debugging.

TUMT. Mitch reconfigured all k2ew modules to be similar to CBHT (i.e. more robust with dropouts). Seems to have fixed TUMT. Now it seems to run out of open connections and only a reboot will fix (since you can't telnet in if there are no connections). Need to look into why sessions don't close.

DBB. Steve suggested a terminal server/watchdog to do dynamic resets when the hang. Has been working great though so this is backburner (need to order a bigger backburner and more pilot lights). ISTI worm module to monitor outages should best wait until RPC is running. Jim will check into getting the dish hung at NMAD and LNXT.

NMAD fence paperwork is ready to install.

GPS. Made several maintenance runs. Really need to get weekly status reports on stations. Need to know what is or isn't communicating before units fill up in the field.

MPHB and the box in the yard is up--sort of. The Lantronix redirector software for making ip:port appear as local serial port is unstable. Working on several avenues. First scream2ew is in beta that will connect to a scream via ip:port instead of just port on the local host. Murray at guralp is working on a new scream that will connect to guralp digitizers over the net instead of just over a local port. ISTI is working on a solaris version of their gcf utils that connect directly to a guralp digitizer via either ip:port or serial, then stream to an earthworm ring.

ET. Mitch shut down the lrnc heli-pages. John is focussing on new installs in east TN. The plan is to completely rebuild the NC stations when we can get time to get over and help. Then we'll put lrnc back online. Receiving all USNSN stations in Memphis. Steve replaced the computers for athn. Now there are no more bsd boxes. Worms running both in the EMA and on star mountain and streaming data. Still need to work out the details of getting data out of athn. Don't have the bandwidth for waveforms, but might send statrig messages and form trigger groups in Memphis.

NM. HALT has a timing problem. Not sure specifically what other stations need work but we're in pretty good shape. Need to work in a routine maintenance trip this summer (e.g. for knocking down the kudzu, checking voltages, etc).

Chapman and Munsey are working on a site for Oakridge to put a CMG3 or 40T. Mitch will provide more info on proposed KY NSN station to get contacts.

FDL. Jim was just figuring out the state of the FDL when we deployed to New York. Need K2 internal batteries and may soon need external batteries. Also need sundry other supplies like plugs and connectors, cables, tools, etc. Could also use some good shipping containers.

NHIN. Heli_ewII on NT apparently has a memory leak. Steve shut it down. Unless the bug is fixed, will have to rely on wave_viewer for the AASG meeting. AASG meeting is June 22 to 26.

PII. Most of PI has been cleaned out. PII proposal for New Madrid was not funded.

Monroe. Still want to. But when and where?

RPC. Parts are being ordered. Sheila set up a constructed equipment account. Mitch is keeping a folder.

FY02 ANSS. Will be installing 20 CMG5TD systems with vaults and accessories as close to last year as possible. See the ANSS page for station details.

Greg is working on version 101f of the I-40 installation.

Need to do site visits of Lamont stations sometime this year.

3904/215E. Will expand into the closet and be used as a computer setup/testing lab.

Generator. Need one.

Purchase CERI SUV. Done.

Purchase Parker Vehicle. Mitch xferred the money. Chris sent the specs to purchasing to initiate a RFB.

Need to organize a Giddings workshop this summer.

ESEE. Postponed to October. Will be 4 CMG40T and 4 L22 with reftek's.

Auburn. Wants to make some cores. Mitch is getting details.